# Playful DoggyBot: Learning Agile and Precise Quadrupedal Locomotion

### v2

@@ IV-A Experiment Setup
We deploy our policy on the Unitree Go2 robot with 12 joints and one extra mouth-like gripper that weighs 15kg. The height of the standing base is about 30cm and its body length is about 70cm. We run the policy on a Jetson Orin NX mounted on the robot. For detecting the target, we use an Intel RealSense D435i mounted on the back of the quadrupedal robot for exteroception, which captures RGB-image at $30$ Hz with an HSV detector to obtain the relative position of the target ball with a diameter of no more than 5cm. Due to the high-speed motion of the robot, the collected images are prone to motion blur, resulting in deviations in the object’s position and corresponding depth information. Therefore, we apply filtering to the received depth information to eliminate abrupt noise. Since our focus is not on hardware and our task requires the gripper to be able to instantly close and grasp objects, we modify a passive gripper with strong magnets that can be swiftly triggered. We also track different objects using Grounding DINO [52] open-vocabulary detector and Mobile SAM [53] tracker. This pipeline works on Nvidia Jetson NX at $8\pm 2$ Hz, although this frequency is not very sufficient for the robot to accurately catch when needing to jump up, it can still show the ability to track objects and play with humans. This indicates that our agile locomotion skills can be integrated with the system based on vision-language models (VLMs), such as in a helping task [4], allowing us to perform a broader range of tasks.
@@ IV-B Simulation Experiments
Figure 4: Comparisons. (a) We examine the success rate of catching the target object while introducing uniform noise to the goal position at various heights. (b) We analyze the success rate of the training range for the target height across three different backbone architectures. (c) and (d) We evaluate the success rate based on whether the input includes the absolute target height, again comparing the performance of the three backbone models.
@@
The robot’s successful attempt is defined as when it jumps, catches the target ball, and then lands without falling over. We test the success rate of the robot at different target heights from 0.5m to 0.9m in the simulation environment, as shown in Table I. We also present keyframe animations of the robot grasping different target objects in Figure 1. Due to the significant impact of the allowable torque on the robot’s jumping ability in the simulation, failing to impose a torque limit could result in a large sim-to-real gap, making the success rates in the simulation less meaningful. Thus we set a maximum torque limit of 30 Nm for the robot in the simulation. We selected the checkpoints of the three best-performing corresponding policy networks for testing. The policy networks based on MLP, GRU, and Transformer achieve success rates of over 80% at a target height of 0.5m, with MLP performing the worst. However, when the target height increases to 0.7m, the success rate decreases as the task difficulty rises, with MLP dropping to around 35%, while GRU and Transformer maintained success rates above 70%. As the target height increased, the MLP struggled due to its lack of memory, resulting in failure at 0.9m, while both the GRU and Transformer achieved success rates exceeding 40%. The overall performance of the GRU and Transformer is quite similar. During the training process, we found that as the required grabbing height increases from a level that did not require jumping to a height that did, a lack of careful adjustment to the reward could result in the model failing to learn that it could jump, leading it to get stuck in a local optimum. The policy based on the Transformer architecture demonstrates some advantages in escaping from this local optimum.
@@
Figure 5: Training Process. When broadening the height range of objects that robots can grasp through curriculum settings, we observe that the success rate for a specific height may decline due to a reduced number of environments associated with that height.
@@
During the training process, we use a curriculum to adjust the height of the target objects, preventing the agent from giving up on catching the ball when faced with excessively difficult targets. When expanding the height range of objects grasped by robots by curriculum setting, we noticed that the success rate of a specific height may decrease due to the decrease in the number of environments corresponding to this height. In one training session shown in Figure 5, we set up 4096 parallel environments and applied height curriculum to dynamically adjust the task difficulty. During the training process, while the number of environments set to a height of 0.5m decreases to 25% of the total number of environments, the success rate at that height exhibits a downward trend. As shown in Figure 4 (b), as the height range increases from 0.1m to 0.5m, the average success rate decreases. We speculate that this may be due to the offset of data distribution or the network capacity not being able to handle the increasing range. However, using different backbones or scaling up the model capacity does not resolve this issue. We then hypothesize that this may also result from the robot’s potential pitch during movement, which makes the representation of the target’s height using relative position less clear. Therefore, we introduce a hint of the target object’s absolute height into the robot’s observation input to compare the impact of this hint on the success rate. Figures 4 (c) and (d) illustrate the two scenarios with and without the absolute height hint. However, the inclusion of the hint significantly impacts the success rate only for the MLP backbone at heights above 0.7m, which may be attributed to MLP’s lack of memory capability. We speculate that this may also be related to the different gaits the robot exhibits when grasping objects at varying heights, as shown in Figure 1. After the curriculum updates, training on a new target may have affected the model’s performance on the original target. This is also a point we can explore further in our future work.
@@ IV-C Real-World Experiments
We quantitatively test our policy in the real world. We select the best-performing checkpoints based on GRU and Transformer backbones from the simulator for testing on the Unitree Go2 real robot. The success rates of different heights of the target ball over 10 trials each are shown in Table I. When the target object’s height is just slightly above the robot’s standing height, the success rate of grasping the object is very high, more than 80%. However, as the height of the object increases, the difficulty of the task significantly rises, leading to a corresponding decrease in the success rate. Due to the significant noise and latency in the relative position information of the target object captured by the camera while the robot is in high-speed motion, the success rate in the real world is lower compared to experiments at the same height conducted in the simulation. When the target height is 0.7m, the robot’s success rate is approximately 30%; however, when the target height increases to 0.8m, the success rate drops to 10%. There exists a sim-to-real gap, as the real robot cannot achieve the highest height successfully reached in the simulation. Our real robot can catch the highest ball at 0.8m, roughly twice the height at which the robot is standing. During the experiments, we found that minimizing the difference in camera latency between the real world and training could improve the success rate of target grabbing. The model that performs best on the real robot has a latency sample range of [0.03, 0.08] seconds during the training process. We test our policy in several different indoor and outdoor environments, including relatively smooth surfaces, grassy areas, and uneven outdoor stone pathways. The robot can catch a fixed ball or play with a moving ball held by a human, as shown in Figure 1. When combined with some more powerful detectors, it can be quite playful.

### v3

@@ IV-A Experiment Setup
We deploy our policy on a Unitree Go2 robot with 12 joints and one extra mouth-like gripper with a total weight of 15kg. The height of the standing base is about 30cm and its body length is about 70cm. We run the policy on a Jetson Orin NX mounted on the robot. For detecting the target, we use an Intel RealSense D435i mounted on the back of the quadrupedal robot for exteroception with a frame rate of 30Hz. Since our focus is not on hardware and our task requires the gripper to be able to instantly close and grasp objects, we design a passive gripper with strong magnets that can be swiftly triggered.
@@ IV-B Simulation Experiments
Figure 4: Comparisons. (a) We examine the success rate of GRU for catching the target object while introducing uniform noise to the goal position at various heights. (b) We analyze the success rate of the training range for the number of different target heights during training across three different backbone architectures. (c) and (d) We evaluate the success rate based on whether the input includes the absolute target height, again comparing the performance of the three backbone models.
@@ IV-B1 Success rate
A robots’s attempt is classified as success if it jumps, catches the target ball, and then lands without falling over. We consider it successful if the robot catches the ball within three attempts of its front legs leaving the ground, as the robot sometimes requires real-time adjustment information to better know the desired jumping height. We test the success rate of the robot at different target heights from 0.5m to 0.9m in the simulation environment, as shown in Table I. We also present keyframe animations of the robot grasping different target objects in Figure 1. Due to the significant impact of the allowable torque on the robot’s jumping ability in the simulation, failing to impose a torque limit could result in a large sim-to-real gap, making the success rates in the simulation less meaningful. Thus, we set a maximum torque limit of 30 Nm for the robot in the simulation. We selected the checkpoints of the three best-performing corresponding policy networks for testing. The policy networks based on MLP, GRU, and Transformer achieve success rates of over 80% at a target height of 0.5m, with MLP performing the worst. However, when the target height increases to 0.7m, the success rate decreases as the task difficulty rises, with MLP dropping to around 35%, while GRU and Transformer maintained success rates above 70%. As the target height further increases, the MLP struggles due to its lack of memory, resulting in failure at 0.9m, while both the GRU and Transformer achieve success rates exceeding 40%. The overall performance of the GRU and Transformer is quite similar. During the training process, we found that as the required grabbing height increases from a level that did not require jumping to a height that did, a lack of careful adjustment of the reward could result in the model failing to learn that it could jump, causing it to get stuck in a local optimum. The policy based on the Transformer architecture demonstrates some advantages in escaping from this local optimum.
@@
Figure 5: Curriculum Learning. When broadening the height range of objects that robots can grasp through curriculum settings, we observe that the success rate for a specific height may decline due to a reduced number of episodes associated with that height.
IV-B2 Curriculum Learning
During the training process, we use a curriculum to adjust the height of the target objects, preventing the agent from giving up on catching the ball when faced with difficult targets. When expanding the height range of objects grasped by robots by curriculum setting, we noticed that the success rate of a specific height may decrease due to the decrease in the number of environments corresponding to this height. In one training session shown in Figure 5, we set up 4096 parallel environments and applied height curriculum to dynamically adjust the task difficulty. During the training process, while the number of environments set to a height of 0.5m decreases to 25% of the total number of environments, the success rate at that height exhibits a downward trend.
@@ IV-B2 Curriculum Learning
As shown in Figure 4 (b), as the number of different heights that the network is trained to master increases, the average success rate goes down. We then hypothesize that this may result from the robot’s potential pitch during movement, which makes the representation of the target’s height using relative position less clear. Therefore, we introduce a hint of the target object’s absolute height into the robot’s observation input to compare the impact of this hint on the success rate. Figures 4 (c) and (d) illustrate the two scenarios with and without the absolute height hint. However, the inclusion of the hint significantly impacts the success rate only for the MLP backbone at heights above 0.7m, which may be attributed to MLP’s lack of memory capability. We speculate that this may also be related to the different gaits the robot exhibits when grasping objects at varying heights, as shown in Figure 1. After the curriculum updates, training on a new target may have affected the model’s performance on the original target. This is also a point we can explore further in our future work.
@@ IV-C Real-World Experiments
We also quantitatively test our policy in the real world. We select the best-performing checkpoints based on GRU and Transformer backbones from the simulator for testing on the Unitree Go2 real robot. The success rates of different heights of the target ball over 10 trials each are shown in Table I. When the target object’s height is 0.5m - just slightly above the robot’s standing height - the success rate of grasping the object is very high, more than 80%. However, as the height of the object increases, the difficulty of the task significantly rises, leading to a corresponding decrease in the success rate.
Due to the significant noise and latency in the relative position information of the target object captured by the camera while the robot is in high-speed motion, the success rate in the real world is lower compared to experiments at the same height conducted in the simulation. When the target height is 0.7m, the robot’s success rate is approximately 30%; however, when the target height increases to 0.8m, the success rate drops to 10%.
There exists a sim-to-real gap, as the real robot cannot achieve the highest height successfully reached in the simulation. The real robot can catch the highest ball at 0.8m, roughly twice the height at which the robot is standing. During the experiments, we found that minimizing the difference in camera latency between the real world and training, by adding a delay to providing the simulated object pose, does improve the success rate of target grabbing. The model that performs best on the real robot has a latency sample range of [0.03, 0.08] seconds during the training process. We test our policy in several different indoor and outdoor environments, including relatively smooth surfaces, grassy areas, and uneven outdoor stone pathways. The robot can catch a fixed ball or play with a moving ball held by a human, demonstrating the ability to continuously follow a mid-air ball moving at speeds of up to 3m/s with direction changes, maintaining a close pursuit without losing track of the target, as shown in Figure 1 and the video.
@@ V CONCLUSION, LIMITATION and FUTURE WORK
We also did some preliminary work to enable the robot pursue and catch arbitrary objects using the Grounding DINO [54] open-vocabulary detector and Mobile SAM [55] tracker. This pipeline works on the Nvidia Jetson NX at $8\pm 2$ Hz. Although this frequency is not very sufficient for the robot to accurately catch when needing to jump up, it shows our ability to track different objects and play with humans. This indicates that our agile locomotion skills can be integrated with the system based on vision-language models (VLMs), such as in a helping task [4], allowing us to perform a broader range of tasks, which we will pursue in the future.